OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (ikram002p):

parallel means both have the same slope so find the slope for a line that passes through (4, 6) and (−1, −4) if u have two points (x,y ) (x0,y0) then the slope =m \(\Huge m=\frac{y-y_0}{x-x_0}\) so the slope for line that passes through (4, 6) and (−1, −4) \(\Huge m=\frac{6--4}{4--2}=\frac{6+4}{4+2}=\frac{10}{6}\)

OpenStudy (anonymous):

so it's (10,6)?

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

no its not , its only 10/6

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (ikram002p):

np , ur wlc :D